Circular No. 6779

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The High-Z Supernova Search Team (IAUC 6160, 6646) reports the

discovery of five supernovae on CCD images taken at the Canada-France-Hawaii Telescope by P. Garnavich, Harvard-Smithsonian Center for Astrophysics (CfA); and H. Dahle, G. Luppino, and D. Clowe, University of Hawaii (UH). The supernovae were identified by B. Schmidt, Mount Stromlo and Siding Spring Observatories; J. Tonry, UH; and P. Challis, CfA, by subtracting images taken during Sept. 26-30 UT from those obtained on Oct. 29–Nov. 1. The limiting magnitude for both the template and search frames was I ∼ 24.5. The supernovae were confirmed on images taken with the WIYN telescope on Nov. 4–5.

SN  1997 UT  α_{2000}  δ_{2000}  I  z type
1997dv Oct. 29 4^{h}15^{m}00.^{s}01 + 5^{o}32′06′′.1 23.0 ?  ?
1997dw Oct. 29 4 14 49.05 + 5 54 51.9 21.2 ?  ?
1997dx Oct. 29 1 38 21.10 −12 43 22.8 22.5 0.21 Ia
1997dy Nov. 1 4 16 35.32 + 5 47 03.1 23.9
1997dz Nov. 1 1 56 03.69 −13 43 06.4 22.3

The supernovae are within 1′′ of their host galaxies except SN 1997dx, which is 2′′.2 west and 0′′.8 north. Spectra of SNe 1997dv, 1997dw, and 1997dx were obtained under poor conditions with the Multiple-Mirror Telescope on Nov. 4–5 by P. Challis and S. Jha (CfA). Finder charts may be requested

by sending electronic mail to pgarnavich@cfa.harvard.edu.

GS 1354–64 R. P. Fender, Astronomical Institute ‘Anton Pannekoek’, University

of Amsterdam; S. J. Tingay, Jet Propulsion Laboratory, NASA; and J. Higdon, R. Wark, and M. Wieringa, Australia Telescope Compact Array (ATCA) Alert Team, report: “We have detected a probable radio counterpart to the blackhole-candidate x-ray transient GS 1354–64, which is undergoing an x-ray and optical outburst (IAUC 6772, 6774, 6775). The radio counterpart lies within 1′′.5 of the optical coordinates. The source has been detected with the ATCA on both Nov. 20/21 and Nov. 24/25 at a level of 1–3 mJy, at frequencies of 2.5, 4.8, and 8.6 GHz. The radio spectrum on Nov. 24/25 corresponds to optically-thin synchrotron emission with a spectral index of ∼ −0.5 and may have evolved from a more absorbed state on Nov. 20/21.”

1997 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green